OpenStudy (anonymous):

Find the LCD that would eliminate the fractions in this equation. 4n-1/7=5/7

OpenStudy (anonymous):

Multiply the whole fraction by seven. You get\[28n-1=5\] Gathering the terms,\[28n=6\] Dividing by 28, \[n=3/14\] I'm not sure what you mean by LCD, but you could, alternatively, Add 1/7 to both sides,\[4n=6/7\] Multiply the left-hand side by 7/7 (you can do that because it's equal to one, so the equality is preserved),\[28n/7=6/7\] Multiply by seven,\[28n=6\] And divide by 28, \[n=6/28\]

OpenStudy (anonymous):

no ur wrong

OpenStudy (anonymous):

Indeed?

OpenStudy (anonymous):

I'm sorry, I should have put 3/14 at the end of the last thing, but really, I hope you can simplify fractions on *that* level on your own.

OpenStudy (anonymous):

i cant im 5

OpenStudy (anonymous):

Right. Well, if you can't figure out that 6/28 = 3/14, maybe these are a bit too high level for you.